OpenStudy (anonymous):

what is x=8+y and x-11y=-12

OpenStudy (anonymous):

This is long

OpenStudy (anonymous):

Here we go Step: Solve x=y+8 for x: Step: Substitute (y+8) for x in x−11y=−12: x−11y=−12 y+8−11y=−12 −10y+8=−12(Simplify both sides of the equation) −10y+8+−8=−12+−8(Add -8 to both sides) −10y=−20 −10y −10 = −20 −10 (Divide both sides by -10) y=2 Step: Substitute (2) for y in x=y+8: x=y+8 x=2+8 x=10(Simplify both sides of the equation) Answer: x=10 and y=2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome